OpenStudy (anonymous):

Put 2x^2+3x=(3x-1)^2-x^2 into ax^2+bx+c=0

OpenStudy (anonymous):

I have tried but I don't understand.

OpenStudy (anonymous):

Start by expanding out the \[(3x-1)^2\] bit.

OpenStudy (anonymous):

I think it'll be easier after that.

OpenStudy (anonymous):

I got 9x^2-3x-3x+1

OpenStudy (anonymous):

-3-3 is -6

OpenStudy (anonymous):

nice dat u've tried. 2x^2+3x=9x^2+1-6x-x^2 =>2x^2-8x^2+3x+6x-1=0 therefore, -6x^2+9x-1=0

OpenStudy (anonymous):

So the answer is 7x^2-9x+1=0?

OpenStudy (anonymous):

\[(3x-1)^2 = 9x^2 -6x +1\] \[\implies (3x-1)^2-x^2 = 8x^2 - 6x + 1\]

OpenStudy (anonymous):

Then use that in your original and you should get the right thing.

OpenStudy (anonymous):

I meant +1

OpenStudy (anonymous):

??? I still don't get it. is it 7x^2-9x+1=0?

OpenStudy (anonymous):

@hpslytherin ... getto see my work out, guess u'll understand.